OpenStudy (anonymous):

to solve : -3c=5c+16 . is it : -3c=5c+16 -(-3c)=-(-3c) 0= 8c +16 16=8c 16/8=8c/8 ???

OpenStudy (anonymous):

c=2

OpenStudy (anonymous):

but my answer sheet says is c = -2

OpenStudy (anonymous):

sorry is c=-2 -3c=5c+16 -8c=16/*(-1) 8c=-16 c=-2